A=1-(B/C) Solve for C.
I multiplied thru by C and got: AC=C-B, and that's where I get stuck. Please help?
Thank you.
Toni

Answer by stanbon(75887)

A=1-(B/C) Solve for C.
----
(B/C) = 1-A
----
Invert both sides:
C/B = 1/(1-A)
---
C = B/(1-A)
---------------------
Using your method:
AC = C - B
---
AC-C = -B
C(A-1) = -B
C = B/(1-A)
====================
Cheers,
Stan H.
=============